Two fixed charges, +1.0 x 10-6 C and -3.0 x 10-6 C, are placed 10 cm apart. (a) Where may a third charge be located so that the net force acting on this charge is zero? (b) Is the equilibrium stable or not?

Please explain and show all formulas to solve this problem! Thanks
Answer #1

let the point is at a distance d left to 10^-6 C charge

at that point electric field due to both the charge must be equal

k q1 / d^2 = k q2 / ( d + 10)^2

1 + 10/d = sqrt ( q2 / q1)

1+ 10/d = sqrt ( 3/1)

d = 13.66 cm

it is unstable equilibrium

=========

do comment in case any doubt, will reply for sure.. Goodluck